# On the correlation between Architectural Smells and Static Analysis Warnings

### v1

@@ Abstract
Background. Software quality assurance is essential during software development and maintenance. Static Analysis Tools (SAT) are widely used for assessing code quality. Architectural smells are becoming more daunting to address and evaluate among quality issues.
Objective. We aim to understand the relationships between Static Analysis Warnings (SAW) and Architectural Smells (AS) to guide developers/maintainers in focusing their effort on SAWs more prone to co-occurring with AS.
Method. We performed an empirical study on 103 Java projects totaling 72 million LOC belonging to projects from a vast set of domains, and 785 SAW detected by four SAT, Checkstyle, Findbugs, PMD, SonarQube, and 4 architectural smells detected by ARCAN tool. We analyzed how SAWs influence AS presence. Finally, we proposed an AS remediation effort prioritization based on SAW severity and SAW proneness to specific ASs.
Results. Our study reveals a moderate correlation between SAWs and ASs. Different combinations of SATs and SAWs significantly affect AS occurrence, with certain SAWs more likely to co-occur with specific ASs. Conversely, 33.79% of SAWs act as ”healthy carriers,” not associated with any ASs.
Conclusion. Practitioners can ignore about a third of SAWs and focus on those most likely to be associated with ASs. Prioritizing AS remediation based on SAW severity or SAW proneness to specific ASs results in effective rankings like those based on AS severity.
@@ 1 Introduction
The risk of service degradation and disruption for critical services is exponentially increasing due to the pervasive nature of computing and the widespread adoption of digital services in all areas of daily life [1]. To create a common definition for quality issues, the literature provides the concept of smells [2], referring to quality issues in code [3]. These smells are usually categorized as code smells, low level, coding issues [4, 2], and, recently, in architectural smells, higher abstract issues that concern the software architecture and the software components [5]. More specifically, architectural smells (AS) are “architectural decisions that negatively impact system internal quality” [6], which can occur simply by “applying a design solution in an inappropriate context, mixing design fragments that have undesirable emergent behaviors, or applying design abstractions at the wrong level of granularity” [6]. High-level quality issues are more difficult to spot given the non-direct, and evident connection, to low-level coding issues [7]. Therefore, recent studies stressed the role of AS in architectural degradation, thus hastening architectural technical debt accumulation  [8, 9, 10].
@@
A popular technique to detect quality issues is to leverage Static Analysis Tools (SAT) [11, 12] which analyze the source code without executing it (thus static analyzers), and produce a report containing all the violations, i.e., Static Analysis Warnings (SAW) that the code present. More specifically, SAWs, refer to code violations with a lower-level granularity, compared to AS. Therefore, SAWs do not consider the software architecture of the system under development, commonly focusing on class or method levels [13]. While multiple studies successfully exploited SATs to remediate common quality issues [13, 14, 3, 15, 16], only a few studies focused on developing SATs to detect ASs [5]. Moreover, some studies have investigated the interrelations between code anomalies [17, 18] or code smells [19] and AS [17, 18]. However, to our knowledge, no previous study has analyzed the correlations between SAW and AS. Furthermore, to ensure the safety and stability of services, during development and maintenance, AS remediation should be prioritized [9, 10]. Hence, the portion of code affected by AS is the primary remediation target to prevent critical quality issues.
@@
Therefore, the need to investigate how practitioners and researchers can exploit SAWs to detect and remediate AS is evident. Hence we designed and performed a large empirical study on the possible correlations between four AS and SAW detected by four tools largely adopted by developers [13].
Hence, this study assesses the correlation between SAW and higher-level AS. We did not consider correlations between defects and SAW, as these have been extensively studied in the literature. Additionally, we investigate the correlation of SAW and AS, and how SAW can be utilized to prioritize AS. Discovering a correlation between AS and SAW would enable the use of existing SAT as an effective tool for detecting issues in software architecture. For practitioners, linking SAW to architectural smells can reduce the need for refactoring, thereby saving both cost and effort [20, 21]. Therefore, we can summarize our main contributions as follows:
@@
  • We performed the first, according to our knowledge, large-scale study, i.e., 112 projects, on the correlation of ASs and SAWs. Other studies have considered correlations with more general architectural problems [18, 17] or the correlation between AS and code smells [19].
  • We investigated the contribution of the most correlated SAWs with ASs.
@@
  • We proposed and evaluated AS remediation effort prioritization via SAWs severity.
@@
Our study focused on 112 Java projects from the Qualitas Corpus dataset (QCD) [22], totaling 72 million lines of code analysis and 66,036 packages, 808,208 classes, and 88,460 interfaces examined across various contexts, including development platforms, multimedia, gaming, middleware, compilers, and testing frameworks. We considered a large set of warnings (1142) and four AS based on dependency issues that can critically influence the quality of a software project and its progressive architecture degradation [23].
@@
Our findings reveal a moderate correlation between SAWs and ASs. Moreover, different combinations of SATs and SAWs significantly influence the occurrence of ASs. Therefore specific SAWs are more prone to co-occur with specific ASs. Conversely, we also identified SAWs acting as “healthy carriers,” i.e., not co-occurring with any ASs. Healthy carriers represent 33.79% of our dataset. Our study shows that practitioners can safely drop a third of the SAW in our context and focus on the most AS-prone to address AS without prior AS severity or discovery knowledge effectively. Our study highlighted that prioritizing AS remediation efforts based on SAW severity, or SAW proneness to specific AS with a limited inspection window, results in rankings similar to the optimal one, i.e., ordered by AS severity.
Therefore, we suggest that SAWs can influence AS’s presence. Prioritizing SAWs based on occurrence probability effectively guides remediation. Future research should improve the detection of AS-prone SAWs and explore deeper connections between SAWs and AS for better understanding and prioritization of remediation efforts.
@@ 2 Background
2.1 Static Analysis Tool
This section introduces the tools we employed in our study and the motivation for their selection. SAT analyzes software without running it by inspecting the source code, thus discovering potential quality issues in the code base [8]. Recently, the ease-to-useness of SATs and the subsequent inclusion in CI practices [24] are boosting their popularity. Among the SAT available, we selected the most widely used according to state of the art in detecting SAW [11, 13] such as Sonarqube, Checkstyle, Findbugs, and PMD. While according to architectural issues, our study focuses on architectural smells (AS). Therefore, we also included ARCAN based on our previous experience [5]. In Table 1, we summarized the characteristics of each of the following SATs.
SAW Checkstyle Findbugs PMD SonarQube
@@ 2.1 Static Analysis Tool
Table 1: Summary of Static Analysis Tools Warning (SAW)
@@
SonarQube [^4] http://www.sonarsource.org/ is one of the most common SATs for detecting code quality issues. SonarQube measures several aspects of the source code, such as several lines of code and code complexity, and verifies code compliance with a specific set of coding rules defined for most common programming languages. Violations of coding rules are reported as issues in SonarQube. The estimated time to remove these issues, i.e., remediation effort, is used to compute the remediation cost. SonarQube analysis encompasses quality aspects such as reliability, maintainability, and security. Reliability rules, also named Bugs, create quality issues that “represent something wrong in the code” and that will soon be reflected in a bug. Code smells are considered “maintainability related issues” in the code that decrease code readability and modifiability. It is important to note that the term “code smells” adopted in SonarQube does not refer to the commonly known code smells term defined by Fowler et al. [4], but to a different set of rules.
@@
On the other hand, we employed ARCAN to detect different AS based on dependency issues by computing different metrics proposed by Martin [25], such as those related to instability issues. We refer to software instability as the inability to make changes without impacting the entire system or a large part.
@@
The tool relies on graph database technology to perform graph queries, which lets on higher scalability during the detection process and management of many different kinds of dependencies. The detailed detection techniques for AS have been described in previous works [5]. Moreover, the tool allows the computation of an architectural debt [26] index based on the number of AS detected in a project and the criticality of the AS, evaluated according to each type of AS. The tool’s results were validated on ten open source systems and two industry projects with a high precision value of 100% in the results and 63% recall [5]. Moreover, the results of ARCAN were validated using practitioners’ feedback on four industry projects [27]. In our study, we have considered the following AS detected by the tool:
  1. 1.
    Unstable Dependency (UD): describes a subsystem (component) that depends on other less stable subsystems than itself [28]. This may cause a ripple effect of changes in the system [5]. Detected in packages.
  2. 2.
    Hub-Like Dependency (HD): This smell arises when an abstraction has (outgoing and ingoing) dependencies on a large number of other abstractions [29]. Detected in classes and packages.
  3. 3.
@@
    Cyclic Dependency (CD): refers to a subsystem (component) that is involved in a chain of relations that break the desirable cyclic nature of a subsystem’s dependency structure. The subsystems involved in a dependency cycle can hardly be released, maintained, or reused in isolation. Detected in classes and packages. The Cyclic Dependency AS is detected according to different shapes [30] as described in [5].
@@
We focused on the above AS based on dependency issues since dependencies are very important in software architecture. Components with many dependencies can be considered more critical and expensive to maintain. Moreover, we defined as “healthy” the instances without AS, i.e., a software entity affected by one or more SAWs but without any of the ASs considered in our study.
@@ 2.2 Effort Aware Metrics
Çarka et al. [31] investigated effort-aware metrics (EAM) to improve defect prioritization. In our study, we take inspiration from the definition of PofB and PopT to create three different prioritization rankers for AS. We then evaluate our ranker in a way that is similar to the interpretation of the PofB results. Therefore, we briefly discuss EAMs; according to Çarka et al. [31], two different EAMs exist: normalized by size or not by size. The most known not-normalized EAM is called PofB [32, 33, 34, 35] PofB, or Proportion of Defects Identified by the top xof the codebase, is a measure derived from a prediction model’s rankings. It signifies the percentage of defective entities discovered when analyzing the initial xof the code. Higher PofB indicates more effective rankings, thereby enhancing testing support. For instance, a PofB10 of 30% implies that 30% of defective entities were detected by analyzing 10% of the codebase according to the method’s ranking.
@@ 3.1 Goal, Research Questions, Metrics, and Hypothesis
RQ1. Are SAWs and AS correlated?
Investigating the correlation between AS and SAW enables us to assess the relationship between high-level architectural issues and low-level code quality problems. We aim to determine whether ASs are correlated with SAWs and the strength of this correlation. To address RQ1, we collected SAW and AS presence per software package. To assess statistical significance, we conjectured one hypothesis ($H_{1}$) as follows:
  • $H_{11}$: There is a statistically significant correlation between the presence of static analyzer warnings and the presence of architectural smells.
@@
  • $H_{01}$: There is no statistically significant correlation between the presence of static analyzer warnings and the presence of architectural smells.
We then tested whether the SAW influenced the presence of ASs computing Spearman’s $\rho$ correlation coefficient [41] between the total number of specific SAWs and ASs, averaged per software package (see Section 3.4).
Moreover, different static analyzer warnings can lead to specific architectural smells. Therefore, assessing whether a specific SAW’s presence is a possible indicator of specific AS presence is essential. Therefore, we ask:
RQ2. Do SAWs influence AS presence?
In the long run, architectural smell can silently deteriorate complex software architecture. Hence, we investigate the characteristics of SAW affecting software projects and leading to AS to identify common traits among projects prone to AS and assess further factors contributing to AS emergence and persistence. To address RQ2 we analyzed the co-occurrence of SAW and ASs. Let $\sigma$ be a SAW and $\alpha$ be an AS, and $\eta$ be a software entity. We state that $\sigma$ co-occur with $\alpha$, if given $\eta$ such that $\sigma$ and $\alpha$ affect contemporary $\eta$. To assess statistically significance, we conjectured two hypotheses ($H_{2}$ and $H_{3}$ ) as follows:
  • $H_{12}$: There is a statistically significant difference in the co-occurrence of pairs of SAW and AS.
  • $H_{13}$: There is a statistically significant difference in the co-occurrence of SAT-specific SAWs’ and AS presence.
@@
  • $H_{02}$: There is no statistically significant difference in the co-occurrence of pairs of SAW and AS.
  • $H_{03}$: There is no statistically significant difference in the co-occurrence of SAT-specific SAWs’ and AS presence.
We computed the co-occurrence of each SAW and AS. For each project package, we counted the number of healthy CDs, UDs, and HLs (and combinations). We tested $H_{2}$ and $H_{3}$ with Wilcoxon Paired Signed Ranked Test (WPT) (see Section 3.4).
Finally, assessing the correlation between SAW and AS and understanding how this relationship influences the presence of AS in software projects merely scratches the surface. Indeed, it falls short, considering practitioners’ limited time for remediation efforts. Therefore, prioritizing these efforts becomes paramount. Thus, we ask:
RQ3. Are SAWs effective for AS remediation effort prioritization?
Effort prioritization is essential to avoid wasting resources on noncritical tasks [42]. No prior study has investigated AS prioritization via SAW severity. According to the “no free lunch theorem”, currently, there is no default severity for code quality checks universally accepted by our community [43]. Therefore, in our study, we investigate exploiting SAW characteristics, such as SonarQube rules severity and cost estimation, the prioritization effort for AS severity according to Arcelli Fontana et al. [5]. We referred to recent empirical studies to assess the accepted severity of AS [44, 45, 46]. More specifically, Sas and Avgeriou [46] propose, on average, an ordinal scale of one to ten, where one is the least severe AS and ten is the highest. Hence, we adopted the average value of CD (5), UD (7), and HL (9). Moreover, regarding SAW severity, we compiled a comprehensive dataset in our previous study [2] focused on SonarQube. Without the risk of generalizability issue [43], we focused this RQ on SonarQube-specific SAWs. To address RQ3 we quantify the probability of a SAW inducing an AS ($\mathcal{P}$) as the relative frequency at which the particular SAW contributed to the occurrence of an AS. To assess statistical significance, we conjectured one hypothesis ($H_{4}$) as follows:
  • $H_{14}$: There is a statistically significant difference in the prioritized effort between the three rankers.
@@
  • $H_{04}$: There is no statistically significant difference in the prioritized effort between the three rankers
We computed $\mathcal{P}$ for each SAT, SAW, and AS. We ranked the SAWs based on their severity [2], $\mathcal{P}$, and AS severity [46]. We tested $H_{4}$ with WPT (see Section 3.4).

### v2

@@ Abstract
Context. Software quality assurance is essential during software development and maintenance. Static Analysis Tools (SAT) are widely used for assessing code quality. Architectural smells are becoming more daunting to address and evaluate among quality issues.
Purpose. We aim to understand the relationships between Static Analysis Warnings (“warnings”) and Architectural Smells (“smell”) to guide developers/maintainers in focusing their effort on warnings more prone to co-occurring with smell.
Methods. We performed an empirical study on 103 Java projects totaling 72 million LOC belonging to projects from a vast set of domains, and 785 warnings were detected by three SAT, Checkstyle, Findbugs, PMD, SonarQube, and 4 architectural smells were detected by the ARCAN tool. We analyzed how warnings influence smell presence. Finally, we proposed an smell remediation effort prioritization based on warning severity and warning proneness to specific smells.
Results. Our study reveals a moderate correlation between warnings and smells. Different combinations of SATs and warnings significantly affect smell occurrence, with certain warnings more likely to co-occur with specific smells. Conversely, 33.79% of warnings are “non-co-occurring” with any of the smells in our dataset. This provides an early indicator for potential architectural concerns before resource-intensive architectural analysis is performed.
Conclusion. Practitioners can ignore about a third of warnings and focus on those most likely to be associated with smells. Prioritizing smell remediation based on warning severity or warning proneness to specific smells results in effective rankings like those based on smell severity. While not a substitute for specialized tools like ARCAN, warning-based prioritization provides a pragmatic bridge between low-level warnings and high-level architectural issues, particularly useful in contexts lacking full architectural visibility.
@@ 1 Introduction
The risk of architectural erosion [98, 71] and disruption for critical services is exponentially increasing due to the pervasive nature of computing and the widespread adoption of digital services in all areas of daily life [21]. To create a common definition for quality issues, the literature provides the concept of smells [44], referring to quality issues in code [89]. These smells are usually categorized as code smells, low-level coding issues [30, 44], and, recently, architectural smells, higher-level abstract issues that concern the software architecture and the software components [3].
@@
A popular technique to detect quality issues is to leverage Static Analysis Tools (SAT) [96, 5] which analyze the source code without executing it (thus static analyzers), and produce a report containing all the violations, i.e., Static Analysis Warnings (from now on “warnings”) that the code presents. More specifically, warnings refer to code violations with a lower-level granularity compared to smells. Therefore, warnings do not consider the software architecture of the system under development, commonly focusing on class or method levels [24].
Despite the large body of work on smells detection techniques using architectural reconstruction, graph-based models, or clustering heuristics [60], these methods are often resource-intensive, demand explicit architectural views, or require specialized tooling that may not be readily integrated into developers’ daily workflows.
In contrast, warnings are pervasive in software engineering practice. They are automatically produced, tightly coupled with development environments, and consulted regularly by practitioners. This makes warnings a pragmatic and low-barrier data source, yet their potential to act smells early indicators of architectural erosion remains largely underexplored.
While multiple studies successfully exploited SATs to remediate common quality issues [24, 69, 89, 48, 2], only a few studies focused on developing SATs to detect smells [3]. Moreover, some studies have investigated the interrelations between code anomalies [54, 53] or code smells [28] and smells [54, 53]. However, to our knowledge, no previous study has analyzed the correlations between warning and smells.
Furthermore, to ensure the safety and stability of services, during development and maintenance, smell remediation should be prioritized [66, 74]. Hence, the portion of code affected by smells is the primary remediation target to prevent critical quality issues.
This work investigates whether warnings, widely available and frequently ignored, can serve as useful signals to identify or prioritize architectural remediation efforts. By doing so, we aim to bridge the gap between commonly available code-level signals and hard-to-detect architectural issues, providing practical insights for both researchers and practitioners.
@@
Therefore, the need to investigate how practitioners and researchers can exploit warnings to detect and remediate smells is evident. Hence, we designed and performed a large empirical study on the possible correlations between four smells and warning detected by three tools largely adopted by developers [24].
Our study assesses the correlation between warning and higher-level smells. We did not consider correlations between defects and warnings, smells, these have been extensively studied in the literature. Specifically, we investigate the correlation between warning and smells and how warning can prioritize smells.
Past studies have struggled to show an obvious relationship between warnings and smells [72]. Nonetheless, discovering a correlation between smells and warnings would enable the use of existing SAT smells as an effective tool for detecting issues in software architecture. Importantly, this approach requires no additional tooling or modeling effort, making it highly attractive in industrial contexts. For practitioners, linking warning to architectural smells can reduce the need for refactoring, thereby saving both cost and effort [17, 16].
Therefore, we can summarize our main contributions as follows:
@@
  • To our knowledge, we performed the first large-scale study, i.e., 103 projects, on the correlation of smells and warnings. Other studies have considered correlations with more general architectural problems [53, 54] or the correlation between smells and code smells [28].
  • We investigated the contribution of the most correlated warnings with smells.
@@
  • We proposed and evaluated smells remediation effort prioritization via warnings severity.
@@
Our study focused on 103 Java projects from the Qualitas Corpus dataset (QCD) [90], totaling 72 million lines of code analysis and 66,036 packages, 808,208 classes, and 88,460 interfaces examined across various contexts, including development platforms, multimedia, gaming, middleware, compilers, and testing frameworks. We considered a large set of warnings (1142) and four smells based on dependency issues that can critically influence the quality of a software project and its progressive architecture erosion [27, 51, 61].
@@
Contrary to past studies, our findings reveal a moderate, yet statistically significant, correlation between warnings and smells. Moreover, different combinations of SATs and warnings significantly influence the occurrence of smells. Therefore, specific warnings are more prone to co-occur with specific smells. Conversely, we also identified that warnings non-co-occurring with any smells represent 33.79% of our dataset. Our study shows that practitioners can safely drop a third of the warnings in our context and focus on the most smell-prone areas to address smells without prior smell severity or discovery knowledge effectively. Our study highlighted that prioritizing smell remediation efforts based on warning severity, or warning proneness to specific smells with a limited inspection window, results in rankings similar to the optimal one, i.e., ordered by smell severity. Therefore, we suggest that warnings can moderately influence the presence of smells. Prioritizing warnings based on occurrence probability effectively guides remediation. Future research should improve the detection of smell-prone warnings and explore deeper connections between warnings and smells for better understanding and prioritization of remediation efforts.
@@
Definitions: In the remainder of this paper, we refer to static analysis warnings as ‘warnings’ and architectural smells as ‘smells’.
@@ 2 Background
This section introduces the static analysis tools used, the types of architectural smells considered, and the effort-aware metrics applied to evaluate prioritization.
2.1 Static Analysis Tools and Warnings
This section introduces the tools we employed in our study and the motivation for their selection. SAT analyzes software without running it by inspecting the source code, thus discovering potential quality issues in the code base [20]. Recently, the readily availability of SATs and the subsequent inclusion in Continuous Improvement (CI) practices [106] are boosting their popularity. Among the available SATs, we selected the most widely used according to state of the art in detecting warning [96, 24] such as smells, Sonarqube, Checkstyle, Findbugs, and PMD. According to architectural issues, our study focuses on architectural smells. Therefore, we also included ARCAN based on our previous experience [3]. In Table 1, we summarize the characteristics of each of the following SATs.
SAT Checkstyle Findbugs PMD SonarQube
@@ 2.1 Static Analysis Tools and Warnings
Table 1: Summary of Static Analysis Tools Warning (warnings)
@@
SonarQube [^4] http://www.sonarsource.org/ is one of the most common SATs for detecting code quality issues. SonarQube measures several aspects of the source code, such smells several lines of code and code complexity, and verifies code compliance with a specific set of coding rules defined for most common programming languages. Violations of coding rules are reported smells issues in SonarQube. The estimated time to remove these issues, i.e., remediation effort, is used to compute the remediation cost. SonarQube analysis encompasses quality aspects such smells reliability, maintainability, and security. Reliability rules, also named Bugs, create quality issues that “represent something wrong in the code” and that will soon be reflected in a bug. Code smells are considered “maintainability-related issues” in the code that decrease code readability and modifiability. It is important to note that the term “code smells” adopted in SonarQube does not refer to the commonly known code smells term defined by Fowler et al. [30], but to a different set of rules.
Tool Detection Capability
# rule # type # severity
Checkstyle 173 14 4
Findbugs 424 9 4
PMD 305 8 5
SonarQube 413 3 5
Table 2: Static Analysis Warnings
2.2 Architectural Smells
@@
On the other hand, we employed ARCAN  [^5] https://www.arcan.tech that detects different smells based on dependency issues by computing different metrics proposed by Martin [56], such smells those related to instability issues. We refer to software instability smells the inability to make changes without impacting the entire system or a large part.
@@
The tool relies on graph database technology to perform graph queries, which enables higher scalability during the detection process and management of many dependencies. The detailed detection techniques for smells have been described in previous works [3]. Moreover, the tool allows the computation of an architectural debt [75] index based on the number of smells detected in a project and the criticality of the smells, evaluated according to each type of smell. The tool’s results were validated on ten open source systems and two industry projects with a high precision value of 100% in the results and 63% recall [3]. Moreover, the results of ARCAN were validated using practitioners’ feedback on four industry projects [58]. A research version of the ARCAN tool has been used for this study, while the latest Arcan commercial version is available at   [^6] https://www.arcan.tech/.
In our study, we have considered the following smells: Unstable Dependency (UD): An Unstable Dependency occurs when a stable component depends on a less stable component.  [57]. This can lead to a ripple effect when the unstable component changes.  [3]. Detected in packages. Example: A stable PaymentProcessor component depending on an unstable BetaFeature.
1class BetaFeature {
2 public void experimentalMethod() { /* Unstable feature */ }
3}
4
5class PaymentProcessor {
6 private BetaFeature feature; // Depends on unstable module
7
8 public PaymentProcessor(BetaFeature feature) {
9 this.feature = feature;
10 }
11}
Hub-Like Dependency (HD): A Hub-Like Dependency occurs when an abstraction has many incoming and outgoing dependencies, making it a critical component that is difficult to modify. [88]. Detected in classes and packages.
Example: A DatabaseHelper class used across multiple services.
1class DatabaseHelper {
2 public void connect() { /* Connect to DB */ }
3 public void executeQuery(String query) { /* Run Query */ }
4}
5
6class UserService {
7 private DatabaseHelper dbHelper = new DatabaseHelper();
8}
9
10class ProductService {
11 private DatabaseHelper dbHelper = new DatabaseHelper();
12}
@@
Cyclic Dependency (CD): A Cyclic Dependency occurs when multiple components depend on each other in a cycle, making independent modifications and releases difficult. The subsystems involved in a dependency cycle can hardly be released, maintained, or reused in isolation. Detected in classes and packages. The Cyclic Dependency smells is detected according to different shapes [1] smells described in [3]. Example: A dependency cycle between A, B, and C.
1class A {
2 private B b;
3}
4
5class B {
6 private C c;
7}
8
9class C {
10 private A a; // Cycle formed: A -> B -> C -> A
11}
@@
We focused on the above smells based on dependency issues since dependencies are very important in software architecture. Components with many dependencies can be considered more critical and expensive to maintain. Moreover, we defined smells “non-co-occurring” the instances without smells, i.e., a software entity affected by one or more warnings but without any of the ASs considered in our study.
@@ 2.3 Effort Aware Metrics
Çarka et al. [12] investigated effort-aware metrics (EAM) to improve defect prioritization. In our study, we take inspiration from the definition of PofB and PopT to create three different prioritization rankers for smells. We then evaluate our ranker in a way that is similar to the interpretation of the PofB results. Therefore, we briefly discuss EAMs; according to Çarka et al. [12], two different EAMs exist: normalized by size or not by size. The most known not-normalized EAM is called PofB [13, 99, 102, 95] PofB, or Proportion of Defects Identified by the top x% of the codebase, is a measure derived from a prediction model’s rankings. It signifies the percentage of defective entities discovered when analyzing the initial x% of the code. Higher PofB indicates more effective rankings, thereby enhancing testing support. For instance, a PofB10 of 30% implies that 30% of defective entities were detected by analyzing 10% of the codebase according to the method’s ranking.
@@ 3.1 Goal, Research Questions, Metrics, and Hypothesis
RQ1. Are warnings and smells correlated?
Investigating the correlation between smells and warnings enables us to assess the relationship between high-level architectural issues and low-level code quality problems. We aim to determine whether smells are correlated with warnings and the strength of this correlation. To address RQ1, we collected warning and smell presence per software package. To assess statistical significance, we conjectured a family of hypotheses, one for each warning-smells pair hypothesis ($H_{1}$) as follows:
  • $H_{11(w,s)}$
    There is a correlation between the number of instances of warning $w$ and the number of instances of architectural smell $s$, averaged at the package level.
@@
  • $H_{01(w,s)}$
    There is no correlation between the number of instances of warning $w$ and the number of instances of architectural smell $s$, averaged at the package level.
Moreover, different static analyzer warnings can lead to specific architectural smells. Therefore, assessing whether a specific warning’s presence is a possible indicator of the presence of specific smells is essential. Therefore, we ask:
RQ2. Do specific types of warnings co-occur with specific types of smells?
In the long run, architectural smell can silently deteriorate complex software architecture. Hence, we investigate the characteristics of warnings affecting software projects and leading to smells to identify common traits among projects prone to smells and assess further factors contributing to the emergence and persistence of smells. To address RQ2, we analyzed the co-occurrence of warnings and smells. Let $\sigma$ be a warning and $\alpha$ be a smell, and $\eta$ be a software entity. We state that $\sigma$ co-occur with $\alpha$, if given $\eta$ such that $\sigma$ and $\alpha$ affect contemporary $\eta$. To assess statistically significance, we conjectured two hypotheses ($H_{2}$ and $H_{3}$ ) as follows:
  • $H_{12{(w,s)}}$
    There is a difference in the co-occurrence of warning $w$ with smell $s$ compared to its co-occurrence with other smells.
  • $H_{13{(w,s)}}$
    There is a difference in the co-occurrence of warning $w$ with smell $s$ compared to its co-occurrence with other smells by SAT.
@@
  • $H_{02{(w,s)}}$
    There is no difference in the co-occurrence of warning $w$ with smell $s$ compared to its co-occurrence with other smells.
  • $H_{03{(w,s)}}$
    There is no difference in the co-occurrence of warning $w$ with smell $s$ compared to its co-occurrence with other smells by SAT.
We computed the co-occurrence of each warning and smell. For each project package, we counted the number of non-co-occurring warnings (NCO) and those affected by CDs, UDs, and HLs (and combinations of them). We tested $H_{2}$ and $H_{3}$ with the Wilcoxon Paired Signed Rank Test (WPT) (Section 3.4).
Finally, assessing the correlation between warnings and smells and understanding how this relationship influences the presence of smells in software projects merely scratches the surface. Indeed, it falls short, considering practitioners’ limited time for remediation efforts. Therefore, prioritizing these efforts becomes paramount. Thus, we ask:
RQ3 Are warnings effective for smell remediation effort prioritization?
Effort prioritization is essential to avoid wasting resources on noncritical tasks [40]. No prior study has investigated smell prioritization via warning severity. According to the “no free lunch theorem”, currently, there is no default severity for code quality checks universally accepted by our community [23]. Therefore, in our study, we investigated exploiting warning and warning attributes detected from SAT, such as smells SonarQube, i.e., rule severity and cost estimation, and the prioritization effort for smell severity according to Arcelli Fontana et al. [3]. We referred to recent empirical studies to assess the accepted severity of smells [29, 72, 81]. More specifically, Sas and Avgeriou [81] proposes, on average, an ordinal scale of one to ten, where one is the least severe smell and ten is the highest. Therefore, we adopt the average smell severity values, CD (5), UD (7), and HL (9), smells proposed by Sas and Avgeriou [81] who empirically derived these ratings based on practitioner feedback and analysis of the impact of smells on maintainability and evolution. Moreover, regarding warning severity, we compiled a comprehensive dataset in our previous study [44] focused on SonarQube. Without the risk of a generalizability issue [23], we focused this RQ on SonarQube-specific warnings. To address RQ3, we quantify the probability of a warning inducing smells ($\mathcal{P}$) smells the relative frequency at which the particular warning contributed to the occurrence of smells. To assess statistical significance, we conjectured one hypothesis ($H_{4}$) as follows:
  • $H_{14}$
    There is a difference in the prioritized effort between the three rankers.
@@
  • $H_{04}$
    There is no difference in the prioritized effort between the three rankers
We computed $\mathcal{P}$ for each SAT, warning, and smell. We ranked the warnings based on their severity [44], $\mathcal{P}$, and smells severity [81]. We tested $H_{4}$ with WPT (see Section 3.4).